# A Radiance Field Loss for Fast and Simple Emissive Surface Reconstruction

### v1

@@ Appendix D Volume relaxation
Figure 19. Equal time PSNR curve plot. Our method converges at a rate comparable to NeRF in the Instant NGP codebase. We have a longer tail in the loss curve since our method spawns more rays per iteration than NeRF.
@@ Appendix E Implementation details
For the geometry reconstruction experiments shown in Figure 12, we used a $L_{1}$ loss to improve convergence in dark regions. Models were trained for $10000$ iterations (reduced from the default $35000$), with the Laplacian weight decaying exponentially to $2\times 10^{-5}$. The Laplacian was estimated via finite differences using six neighboring samples with an epsilon of $1/1024$ (approximately 1 mm for a unit cube).
@@
Ours 0.673 0.918 0.872 0.577 0.686 0.866 0.896 0.769 0.692
Ours (relaxed) 0.682 0.927 0.882 0.590 0.695 0.878 0.902 0.784 0.698
NeRF 0.675 0.924 0.877 0.586 0.687 0.877 0.893 0.776 0.692

### v2

@@ Appendix D Volume relaxation
Figure 19. Equal time convergence plot. Our method converges at a rate comparable to NeRF in the Instant NGP codebase. We have a longer tail in the loss curve since our method spawns more rays per iteration than NeRF.
@@ Appendix E Implementation details
For the geometry reconstruction experiments shown in Figure 12 (main document), we used a $L_{1}$ loss to improve convergence in dark regions. Models were trained for $10000$ iterations (reduced from the default $35000$), with the Laplacian weight decaying exponentially to $2\times 10^{-5}$. The Laplacian was estimated via finite differences using six neighboring samples with an epsilon of $1/1024$ (approximately 1 mm for a unit cube).
@@
Ours 0.673 0.918 0.872 0.686 0.866 0.896 0.769 0.577 0.692
Ours (relaxed) 0.682 0.927 0.882 0.695 0.878 0.902 0.784 0.590 0.698
NeRF 0.675 0.924 0.877 0.687 0.877 0.893 0.776 0.586 0.692
Table 5. LPIPS comparison using the Instant NGP codebase.
Bicycle Bonsai Counter Garden Kitchen Room Stump Flowers Treehill
Ours 0.578 0.241 0.315 0.547 0.236 0.306 0.475 0.618 0.599
Ours (relaxed) 0.642 0.244 0.335 0.672 0.234 0.324 0.497 0.676 0.645
NeRF 0.658 0.256 0.354 0.625 0.239 0.362 0.514 0.699 0.692
@@
Table 7. Chamfer Distance comparison on the DTU dataset with NeuS [Wang et al., 2021] and NeuS2 [Wang et al., 2023].
Scan24 Scan37 Scan40 Scan55 Scan63 Scan65 Scan69 Scan83
Ours (1 minute) 0.81 0.77 0.66 0.40 1.08 0.90 0.88 1.42
NeuS (8 hours) 0.83 0.98 0.56 0.37 1.13 0.59 0.60 1.45
NeuS2 (5 minutes) 0.56 0.76 0.49 0.37 0.92 0.71 0.76 1.22
Scan97 Scan105 Scan106 Scan110 Scan114 Scan118 Scan122
Ours 1.20 0.75 0.68 1.07 0.61 0.55 0.63
NeuS 0.95 0.78 0.52 1.43 0.36 0.45 0.45
NeuS2 1.08 0.63 0.59 0.89 0.40 0.48 0.55